OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2 solve for g(f(3)).

OpenStudy (gorv):

f(3)=2*3+1=6+1=7

OpenStudy (gorv):

g(f(3))=g(7)=(3*7-1)/2=(21-1)/2=20/2=10

OpenStudy (gorv):

@Imbueds

OpenStudy (anonymous):

Thank you Gorv!!!